OpenStudy (anonymous):

A city was founded at the beginning of 1990 with a population of 55,000, and since then, the growth in its population has been exponential, increasing at x percent per year. If the city's population at the beginning of 2000 was 108,193, at what percent per year to the nearest percent is the city's population increasing?

OpenStudy (anonymous):

he basic equationnfor population growth is P(t) = Po*(1+r)^t, where Po is the initial population, t is the number of years, and r is the annual rate. So we have everything but r: 108193 = 55000*(1+r)^10 (1+r)^10 = 108193/55000 Just find the 10th root of 10893/55000 and subtract 1 to get the rate. If you have a scientific calculator that will do x^y, then just punch in (10893/55000)^(0.1) -1

OpenStudy (anonymous):

thank you so much. :)

OpenStudy (anonymous):

he he